OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x 2 sin x cos^2x + cos x 2 sin x cos^2x + sin^3x sin x cos^2x - sin^3x + cos^3x 2 cos^2x sin x + sin x - 2 sin^3x

OpenStudy (anonymous):

(sinx)(cos2x)+(cosx)(sin2x)

OpenStudy (anonymous):

(sinx)[(cosx)^2 - (sinx)^2] + (cosx)[2(sinx)(cosx)]

OpenStudy (anonymous):

@Data_LG2 how do I go from here?

OpenStudy (anonymous):

check this, it's the same thing: http://openstudy.com/study#/updates/50f3105ee4b0694eaccf7977

OpenStudy (anonymous):

None of the answers are applicable

OpenStudy (anonymous):

@FutureMathProfessor I edited the question to make the answers applicable

OpenStudy (anonymous):

shouldn't the first option be 2 sin x cos^2x + sin x ?

OpenStudy (anonymous):

nah... forget about me...

OpenStudy (anonymous):

the fourth option is correct

OpenStudy (anonymous):

How do you get that?

OpenStudy (anonymous):

since \[\sin3x=\sin(x+2x)\]we use double angle formula to get\[\sin(x+2x) = sinxcos2x+cosxsin2x\]using the double angle formula again\[sinx(\cos^{2}x-\sin^{2}x) + cosx(2sinxcosx)\]since \[\cos^{2}x - \sin^{2}x = 1-2\sin^{2}x\]then ,\[sinx(1-2\sin^{2}x)+2sinxcos^{2}x\]expanding out the bracket will give you\[sinx-2\sin^{3}x+2sinxcos^{2}x\]

OpenStudy (anonymous):

Do you understand it?

OpenStudy (anonymous):

Thank you, I was using the wrong identity the whole time

OpenStudy (anonymous):

You really helped me out a lot

OpenStudy (anonymous):

haha... good luck with these kinda questions